When I add 3 times my age 3 years from now to 4 times my age 4 years from now, I get 8 times my current age. How old will I be 2 years from now?"
Answer by lwsshak3(11628)

When I add 3 times my age 3 years from now to 4 times my age 4 years from now, I get 8 times my current age. How old will I be 2 years from now?"
**
let x=present age
3(x+3)+4(x+4)=8x
3x+9+4x+16=8x
x=25
2years from now:
25+2=27
How old will I be 2 years from now? 27